OpenStudy (anonymous):

what is the equation in point-slope form of the line passing through (2, 5) and (–1, 8) ?

OpenStudy (anonymous):

point slope form equation is as follows y - y1 = m(x - x1) m is the slope and x1 and y1 are the coordinates of a point on the line. the slope of any given two points is (y2 - y1)/(x2 - x1) so the slope in this case is (8 - 5)/(-1 - 2) = 3/-3 = -1

OpenStudy (katrinakaif):

First find the slope of the two given points: (2,5) = (x1,y1) (-1,8) = (x2.y2) Use the equation: y2-y1 ---- To find the slope x2-x1 Then plug the slope as "m" into the following formula and using any pair of coordinate points, plug it into the formula as (x1,y1) (y-y1) = m (x-x1) Solve as a linear function and you shall end with a point-slope equation.